# Semi-Blind Strategies for MMSE Channel Estimation Utilizing Generative Priors

### v1

@@ VI Numerical Simulations
Figure 3: NMSE over the SNR for given channel estimations based on $N=200$ observations including one pilot per user in a $J=8$ user scenario using (a) the spatial channel model and (b) measurement data.
To evaluate our proposed methods, we use channel realizations, which are normalized with $\mathbb{E}\left[\|{\bm{h}}\|^{2}\right]=M$. Thus, we can define the SNR $=\frac{1}{\sigma^{2}}$. Further, the normalized MSE (NMSE) defined as
@@
is used to characterize the performance of the estimators based on $L=10^{3}$ unseen channel samples stemming from the channel models detailed in Sections II-A and II-B. The assumption of spatial uncorrelated channels, as used in (25), only holds for the spatial channel model of Section II-A. For the case of the measurement campaign described in Section II-B, we approximate the noise covariance matrix (24) as
@@
We use =$\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ with $T=1.5\cdot 10^{5}$ training samples from the respective channel model to train the GMM and VAE, where we set the number of components to $K=64$ and the latent dimension to $Z=32$, respectively. Further, in the case of the VAE we allow non-zero values for $\bm{\mu}_{\bm{\theta}}({\bm{z}})$, as we use the circulant approximation for the spatial channel model and the block-Toeplitz property is not perfectly fulfilled for the measurement data, due to hardware imperfections. The “s-cov” variants (“sub. s-cov” and “proj. s-cov”) utilize the same training samples. The number of BS antenna is set to $M=64$, cf. Sections II-A and II-B, serving $J=8=M/8$ number of users, a representative operating point [37, Chap. 1.3.3]. Further, the number of snapshots is set to $N=200$, if not stated otherwise, corresponding to a scenario that allows high channel dispersion and high mobility, e.g., up to $135$ kph, c.f. [37, Chap. 2.1]. The sent symbols during data transmission generally stem from a discrete constellation, e.g., QPSK, $16$-QAM. For this work, we utilize Gaussian symbols with $x_{j}(n)\sim\mathcal{N}_{\mathbb{C}}(0,P_{j}=1/J)$ such that $\sum_{j=1}^{J}P_{j}=1$. Using a continuous symbol constellation has a negligible effect on the results of the simulations, as also previously observed in [16].
Fig. 3a and Fig. 3b show the performance of the different channel estimation methods with respect to the SNR for the spatial channel model (cf. Section II-A) and measurement data (cf. Section II-B), respectively. One can see that the semi-blind methods utilizing the CGLMs perform the best across the whole SNR. The projected variant slightly outperforms its subspace counterpart for most SNR values, which follows the derivations in Section III. Interestingly, the order of the semi-blind GMM and semi-blind VAE depends on the utilized channel model. In Fig. 3a, the projected GMM and VAE show both the best overall result, whereas in Fig. 3b, the projected GMM outperforms all other estimators. Additionally, we see that in Fig. 3a, the subspace VAE outperforms the subspace GMM and in Fig. 3b, the results are vice versa. This ordering follows the ordering of the standalone version, where the plain GMM is better than the plain VAE in the case of the measurement data and worse for the spatial channel model. For high SNR values, all semi-blind variants approach each other except the EM and MP methods. In Fig. 3a, the EM and MP drastically improve from $15$dB to $20$dB showing similar performance at $20$dB as the other semi-blind methods, whereas, in Fig. 3b, they show inferior results also for high SNR. The CGLM-based approaches keep a slight advantage even for high SNR values, which can be attributed to the fact that prior information is beneficial even for high SNR. A notable observation is that, in the mid-SNR range, the semi-blind CGLM variants outperform all related estimators by roughly $3$ dB.
@@
Figure 4: NMSE over the number of observations for given channel estimations based on $N$ observations including one pilot per user at SNR $=0$ dB in a $J=8$ user scenario using (a) the spatial channel model and (b) measurement data.
For our proposed strategies, the accuracy of the estimated subspace $\mathrm{range}(\hat{{\bm{V}}})$ influences the performance and, hence, the NMSE depends on the number of snapshots $N$ as shown in Fig. 4. We see that for an increasing number of snapshots, the NMSE of our proposed methods decreases. In the case of the spatial channel model (Fig. 4a) the projected GMM and VAE perform best for low numbers of snapshots, where the standalone VAE surpasses all other methods for $N=20$. Additionally, we observe in Fig. 4a that for high $N$, the subspace VAE becomes the best of all considered methods. In Fig. 4b, we observe again that for the measurement data, the semi-blind GMM variants perform the best, where for high $N$ the subspace GMM and low $N$ the projected GMM outperforms all other methods. Again, for less than $30$ snapshots, the semi-blind methods are outperformed by the standalone GMM and VAE due to inaccuracies in estimating the subspace with a low number of payload data symbols. For both utilized channel models, the subspace variant of the superior CGLM surpasses the projected variant for high numbers of snapshots converging to a lower error level. Thus, in practice, where, in general, uncorrelated Rayleigh fading is not the case, there are cases where the subspace CGLM outperforms its projected counterpart.
A critical decrease in performance can be observed for the EM and MP algorithms. Here, the NMSE increases after a certain point when increasing the number of snapshots. Even though the minimum appears at different $N$, the overall behavior exhibits similarities. This is because both methods optimize the joint ML formulation in (13), where the optimization of the second term becomes dominant for a high number of snapshots. Hence, the impact of the pilot observations relevant to estimating the phase of the channel vanishes.
@@
Figure 5: NMSE over the number of users for given channel estimations based on $N=200$ observations including one pilot per user at SNR $=0$ dB using (a) the spatial channel model and (b) measurement data.
The dimension of the subspace $\mathrm{range}(\hat{{\bm{V}}})$ directly influences the estimation quality of the proposed methods as shown in Fig. 5a and Fig. 5b. For example, in the extreme case where the number of users in the system is equal to the number of BS antennas ($J=M$), the solution to (56) becomes $\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}=\mathbf{I}$ and, hence, as the number of users in the system increases all semi-blind estimators approach their respective purely pilot based version. We restrict our simulations within the interval of $J\leq M/4=16$, which is said to be the preferred operating regime in massive MIMO [37, Chap. 1.3.3], and set the number of snapshots to $N=200$. In the case of a single user, all semi-blind variants exhibit similar performance, except for the subspace sample covariance estimator, the subspace VAE, and in the case of the spatial channel model (Fig. 5a) the subspace GMM. For all other considered numbers of users, the proposed projected CGLM methods outperform all other channel estimators. Additionally, for the spatial channel model in Fig. 5a, the subspace GMM also shows inferior results to the other CGLM-based methods for all numbers of users.
Overall, we can conclude that the proposed semi-blind CGLMs show superior channel estimation performance across all different setups. Depending on the used channel model, either the semi-blind GMMs or the semi-blind VAEs result in slightly better NMSE, where only in the case of the spatial channel model, the subspace GMM shows slightly worse performance compared to the other proposed methods. Moreover, the projected CGLMs outperform their respective subspace counterpart for most simulated operating points, showing the superiority of the proposed projection method.
@@ VII Conclusion
This work presented a novel semi-blind channel estimation technique based on the class of CGLMs. To this end, two methods are discussed that incorporate subspace knowledge about the channel into the well-known LMMSE estimator. Both methods exploit the estimated subspace derived from the dominant eigenvectors of sample covariance matrices constructed using the received symbols. A theoretical analysis of the methods showed the superior estimation quality of the proposed projection-based estimator for uncorrelated Rayleigh fading channels. Secondly, we showed how two examples from the class of CGLMs, i.e., the GMM and VAE, can be used to parameterize these estimators. Extensive simulations based on real-world measurement and spatial channel model data demonstrated the superior estimation performance of the proposed methods compared to standard semi-blind channel estimators.
@@
$=\mathbb{E}\left[\mathrm{tr}({\bm{h}}{\bm{h}}^{\mathrm{H}})-2\mathrm{tr}({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}})+\mathrm{tr}({\bm{W}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})\right].$ (64)
@@
×$\quad\quad\quad\times\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\bigg{)}\bigg{]}$ (71)
@@
where we assume that $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\sigma^{2}\frac{J}{M}\mathbf{I}_{M}$. From this the overall MSE in (28) follows directly.
@@
In the case of uncorrelated Rayleigh fading the subspace LMMSE filter is given as
@@
The overall MSE of the subspace variant for ${\bm{C}}=\mathbf{I}_{M}$ is then

### v2

@@ VI Numerical Simulations
Figure 4: NMSE over the SNR for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots and Gaussian symbols, using (a) the spatial channel model and (b) measurement data.
Figure 5: NMSE over the number of antennas for given channel estimations in a $J=M/8$ user scenario based on $N=25J$ observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using the spatial channel model.
To evaluate our proposed methods, we use channel realizations, which are normalized with $\mathbb{E}\left[\|{\bm{h}}\|^{2}\right]=M$. Thus, we can define the SNR as $\sigma^{-2}$. Further, the normalized MSE (MSE) defined as
@@
is used to characterize the estimators’ performances based on $L=10^{3}$ unseen channel samples stemming from the channel models detailed in Sections II-A and II-B. We use $\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ with $T=1.5\cdot 10^{5}$ training samples from the respective channel model to train the GMM and VAE, where we set the number of components to $K=64$ and the latent dimension to $Z=32$, respectively. Further, in the case of the VAE we allow non-zero values for $\bm{\mu}_{\bm{\theta}}({\bm{z}})$, as we use the circulant approximation for the spatial channel model and the block-Toeplitz property is not perfectly fulfilled for the measurement data, due to hardware imperfections. The VAE’s remaining implementation details are the same as provided in [32]. The assumption of spatial uncorrelated channels, as used in (23), only holds for the spatial channel model of Section II-A. For the case of the measurement campaign described in Section II-B, we approximate the noise covariance matrix in (22) as
@@
Using the $T$ training samples to empirically approximate the projected noise covariance matrix (22) did not result in a different performance compared to using the approximation (64). The “s-cov” variants (“sub. s-cov” and “proj. s-cov”) utilize the same training samples. For most of the simulations, the number of BS antennas is set to $M=64$, cf. Sections II-A and II-B, serving $J=8=M/8$ number of users, a representative operating point [38, Chap. 1.3.3]. Further, if not stated otherwise, the number of snapshots is set to $N=200$, corresponding to a scenario that allows high channel dispersion and mobility, e.g., up to $135$ kph, c.f. [38, Chap. 2.1].
In Fig. 3, the channel estimation performances are compared for different symbol constellations using the proposed GMM based semi-blind channel estimation variants. Additionally, we compare to the purely pilot-based GMM channel estimator with $N_{p}=J$ and $N_{p}=N$ number of pilots, where the latter assumes all sent symbols to be known by the receiver. We utilize Gaussian symbols with $x_{j}(n)\sim\mathcal{N}_{\mathbb{C}}(0,P_{j}=1/J)$ such that $\sum_{j=1}^{J}P_{j}=1$ as well as QPSK and 1616QAM. In the latter two cases, also a decision directed GMM is evaluated, which decodes the sent symbols based on a pilot-based GMM channel estimate and uses the decoded symbols as additional pilots in a second stage. For the decoding, we utilize LMMSE equalization with a subsequent mapping to the closed constellation point.
Firstly, we see no qualitative difference between the different symbol constellations. Using a continuous symbol constellation has a negligible effect on the simulation results, as also previously observed in [15] and, hence, for the rest of this work, we utilize Gaussian symbols. Further, we see in Fig. 3 an improvement from the pilot-only GMM to our proposed semi-blind variants, showcasing again the advantage of utilizing the payload symbols in channel estimation. The full pilot GMM using $N_{p}=N$ pilot symbols achieves the SNR improvement as shown in Section II. The decision directed GMM based estimator saturates for high SNR values, making it an inferior semi-blind strategy.
@@
Figure 6: NMSE over the number of observations for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on NN observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using (a) the spatial channel model and (b) measurement data.
Fig. 4a and Fig. 4b show the different channel estimators’ performances with respect to the SNR for the spatial channel model (cf. Section II-A) and measurement data (cf. Section II-B), respectively. One can see that the semi-blind methods utilizing the CGLMs perform the best across the whole SNR. The projected variants of the CGLMs slightly outperform their subspace counterpart for most SNR values, which follows the derivations in Section III. Interestingly, the order of the semi-blind GMM and semi-blind VAE depends on the utilized channel model. In Fig. 4a, the projected GMM and VAE show both the best overall result, whereas in Fig. 4b, the projected GMM outperforms the VAE-based estimator. Additionally, we see that in Fig. 4a, the subspace VAE outperforms the subspace GMM, and in Fig. 4b, the results are vice versa. This ordering follows the ordering of the standalone version, where the plain GMM is better than the plain VAE in the case of the measurement data and worse for the spatial channel model. This comes from the fact that the Toeplitz assumption at the VAE may not be fulfilled due to impairments in the real measurement data. For high SNR values, all semi-blind variants approach each other except the EM and MP methods. In Fig. 4a, the EM and MP drastically improve from 1515dB to 2020dB, showing similar performance at 2020dB as the other semi-blind methods, whereas, in Fig. 4b, they show inferior results also for high SNR. The CGLM-based approaches keep a slight advantage even for high SNR values, which can be attributed to the fact that prior information is beneficial even for high SNR. A notable observation is that, in the mid-SNR range, the semi-blind CGLM variants outperform all related estimators by roughly $3$ dB.
In Fig. 5 we evaluate the respective estimators across different numbers of antennas at the BS. In order to make the comparison fair, we keep the number of users at a constant ratio of $J=M/8$ and set the number of snapshots to $N=25J$. With this, we only see an improvement for higher numbers of antennas for the CGLMs, while the other methods roughly keep their performance. Due to the constant parameter ratio, the VAE variants’ drastic performance gains for a higher number of antennas solely come from the VAE’s strong capability of modeling high-dimensional data.
[…]
A critical decrease in performance can be observed for the EM and MP algorithms. Here, the NMSE increases after a certain point when increasing the number of snapshots. Even though the minimum appears at different $N$, the overall behavior exhibits similarities. This is because both methods optimize the joint ML formulation in (13), where the optimization of the second term becomes dominant for a high number of snapshots. Hence, the pilot observation’s impact, which is relevant to estimating the channel phase, vanishes.
@@
Figure 7: NMSE over the number of users for given channel estimations in a $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using (a) the spatial channel model and (b) measurement data.
[…]
Overall, we can conclude that the proposed semi-blind CGLMs show superior channel estimation performance across all different setups. Depending on the used channel model, either the semi-blind GMMs or the semi-blind VAEs result in slightly better NMSE, where only in the case of the spatial channel model, the subspace GMM shows slightly worse performance compared to the other proposed methods. Moreover, the projected CGLMs outperform their respective subspace counterpart for most simulated operating points, showing the proposed projection method’s superiority.
@@ VII Conclusion
This work presented a novel semi-blind channel estimation technique based on the class of CGLMs. To this end, two methods that incorporate subspace knowledge into the well-known LMMSE estimator are discussed. Both methods exploit the estimated subspace derived from the dominant eigenvectors of sample covariance matrices constructed using the received symbols. A theoretical analysis of the methods showed the proposed projection-based estimator’s superior estimation quality for uncorrelated Rayleigh fading channels. Secondly, we showed how two examples from the class of CGLMs, i.e., the GMM and VAE, can be used to parameterize these estimators. Extensive simulations based on real-world measurement and spatial channel model data demonstrated the proposed methods’ superior estimation performances compared to standard semi-blind channel estimators.
@@
$=\mathbb{E}\left[\mathrm{tr}({\bm{h}}{\bm{h}}^{\mathrm{H}})-2\mathrm{tr}({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})+\mathrm{tr}({\bm{W}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})\right].$ (66)
@@
$\quad\quad\quad\times\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\bigg)\bigg]$ (73)
@@
where we assume that $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\sigma^{2}\frac{J}{M}\mathbf{I}_{M}$. From this, the overall MSE in (26) follows directly.
@@
In the case of uncorrelated Rayleigh fading, the subspace LMMSE filter is given as
@@
The overall subspace variant’s MSE for ${\bm{C}}=\mathbf{I}_{M}$ is then
@@
-C Derivation of Bayesian Cramer-Rao Bounds
The BCRB, which is the lower bound on the MSE of any estimator $\hat{{\bm{h}}}({\bm{Y}})$, is given as [48]
$\mathbb{E}\left[\|\hat{{\bm{h}}}({\bm{Y}})-{\bm{h}}\|^{2}\right]\geq\mathrm{tr}\left(\left[\left(\mathbb{E}_{{{\bm{h}}}}[\mathcal{I}({{{\bm{h}}}})]+\bm{J}_{p}\right)^{-1}\right]_{:M,:M}\right),$ (88)
where $\mathcal{I}({{{\bm{h}}}})$ is the Fisher information matrix corresponding to the inverse of the unbiased CRB, and
$\bm{J}_{p}$ $=\mathbb{E}_{\bm{h}}\left[\frac{\partial\log p({\bm{h}})}{\partial{\bm{h}}^{\mathrm{H}}}\frac{\partial\log p({\bm{h}})}{\partial{\bm{h}}}\right]$ (89)
$=\mathrm{blckdiag}({{\bm{C}}}_{1},\dots,{{\bm{C}}}_{J})^{-1},$ (90)
We denote with $\cdot]_{:M,:M}$ in (88) the part of the overall BCRB matrix corresponding to the upper left block of size $M\times M$, where we assume without loss of generality to index the users such that the user of interest is the first one.
For the case of ${\bm{C}}_{j}=\mathbf{I}_{M}$ for all $j\in\{1,\dots,J\}$ the prior information $\bm{J}_{p}$ is given as
$\bm{J}_{p}=\mathbf{I}_{JM}.$ (91)
For the deterministic BCRB, we use the formulation derived in [40], which results in
$\mathbf{BCRB}_{\mathrm{iid,d}}$
$\quad=\sigma^{2}\left(({\bm{X}}^{*}{\bm{X}}^{\mathrm{T}})\otimes{\bm{P}}_{\bm{H}}^{\perp}+({\bm{P}}^{*}{\bm{P}}^{\mathrm{T}})\otimes{\bm{P}}_{\bm{H}}+\sigma^{2}\mathbf{I}_{JM}\right)^{-1},$ (92)
with ${\bm{P}}_{\bm{H}}^{\perp}=\mathbf{I}_{M}-{\bm{P}}_{\bm{H}}$.
Now using the Fisher information matrix for the stochastic CRB as provided in [15], we introduce $\bar{{\bm{h}}}=[\Re\{{\bm{h}}_{1}^{\mathrm{T}}\},\Im\{{\bm{h}}_{1}^{\mathrm{T}}\},\dots,\Re\{{\bm{h}}_{J}^{\mathrm{T}}\},\Im\{{\bm{h}}_{J}^{\mathrm{T}}\}]^{\mathrm{T}}$, which gives us the prior information as
$\bm{J}_{p}$ $=\mathbb{E}_{\bar{{\bm{h}}}}\left[\frac{\partial\log p(\bar{{\bm{h}}})}{\partial\bar{{\bm{h}}}^{\mathrm{T}}}\frac{\partial\log p(\bar{{\bm{h}}})}{\partial\bar{{\bm{h}}}}\right]$ (93)
$=2\,\mathrm{blckdiag}(\bar{{\bm{C}}}_{1},\dots,\bar{{\bm{C}}}_{J})^{-1},$ (94)
with
$\bar{{\bm{C}}}_{j}=\begin{bmatrix}\Re\{{\bm{C}}_{j}\}&-\Im\{{\bm{C}}_{j}\}\\ \Im\{{\bm{C}}_{j}\}&\Re\{{\bm{C}}_{j}\}\end{bmatrix}$. (95)
Thus, the BCRB for uncorrelated Rayleigh fading is given as
$\mathbf{BCRB}_{\mathrm{iid,s}}=\left(\left(\frac{2N_{p}}{J\sigma^{2}}+2\right)\mathbf{I}_{2JM}+{\bm{R}}\right)^{-1},$ (96)
where $\bm{R}$} denotes the part depending on the payload symbols defined in [15].
-D Proof of Theorem 2
To this end, let us consider the MSE for a fixed channel matrix $\bm{H}$}, where the expectation is only taken over the noise and Gaussian symbol realizations as
$\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]$
$\quad=\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|{\bm{W}}_{\mathrm{proj}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{y}}\|^{2}]$ (97)
$\quad\leq\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})\mathbb{E}_{{\bm{X}}}\left[(\mathrm{tr}({\bm{A}})+\mathbb{E}_{{\bm{N}}}[\mathrm{tr}({\bm{B}})])\right]$ (98)
with
$\mathrm{tr}({\bm{A}})$ $=\mathrm{tr}((\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{h}}{\bm{h}}^{\mathrm{H}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}))$ (99)
$=\mathrm{tr}((\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{h}}{\bm{h}}^{\mathrm{H}}(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}))$ (100)
$=\|(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{h}}\|^{2}$ (101)
$=\|(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{V}}{\bm{s}}\|^{2}$ (102)
$=\sum_{i\leq J}\sum_{j>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}s_{i}|^{2}$ (103)
$\leq|s_{\max}|^{2}\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (104)
$=|s_{\max}|^{2}\sum_{i\leq J}\sum_{j>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (105)
where ${\bm{h}}={\bm{V}}{\bm{s}}$ with ${\bm{s}}=[s_{1},\dots,s_{J}]^{\mathrm{T}}$ is the channel of interest’s decomposition in terms of $\bm{V}$}, and $s_{\max}=\max_{i}s_{i}$. Furthermore, we have
$\mathbb{E}_{\bm{N}}[\mathrm{tr}({\bm{B}})]$ $=\mathbb{E}_{\bm{N}}[\mathrm{tr}((\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{n}}{\bm{n}}^{\mathrm{H}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}))]$ (106)
$=\sigma^{2}(\mathrm{tr}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}+{\bm{V}}{\bm{V}}^{\mathrm{H}})-2\mathrm{tr}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}))$ (107)
$=2\sigma^{2}\left[J-\sum_{i,j\leq J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}\right]$ (108)
$=2\sigma^{2}\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (109)
Thus,
$\mathbb{E}_{{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]$
$\quad\leq\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})\left(|s_{\max}|^{2}+2\sigma^{2}\right)\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}.$ (110)
Now, let us consider the following corollary from [49, Corollary 4.1] about the inner product of eigenvectors of sample and true covariance matrices.
Corollary 2.1.
Let us denote with $\lambda^{\prime}_{j}$ and ${\bm{v}}_{j}$ the $i$-th eigenvalue and eigenvector of the true covariance matrix, respectively. Then, for any weights $w_{ji}$, which are non-zero when $\lambda^{\prime}_{i}\neq\lambda^{\prime}_{j}$ and $\mathrm{sgn}(\lambda^{\prime}_{i}-\lambda^{\prime}_{j})2\lambda^{\prime}_{i}>\mathrm{sgn}(\lambda^{\prime}_{i}-\lambda^{\prime}_{j})(\lambda^{\prime}_{i}+\lambda^{\prime}_{j})$, and real $\varepsilon>0$
$P\left(\sum_{j\leq J}\sum_{i>J}w_{ji}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}\leq\varepsilon\right)$
$\quad\geq 1-\sum_{j\leq J}\sum_{i>J}\frac{4k^{2}_{i}w_{ji}}{N\varepsilon(\lambda^{\prime}_{j}-\lambda^{\prime}_{i})^{2}},$ (111)
where $k_{i}^{2}=\mathbb{E}[\|{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{v}}_{i}\|^{2}]-\lambda_{i}^{\prime,2}$.
In our given problem, the true covariance matrix is given as ${\bm{C}}_{{\bm{y}}\mid{\bm{H}}}={\bm{H}}{\bm{H}}^{\mathrm{H}}/J+\sigma^{2}\mathbf{I}_{M}$. It is then trivial to see that $\lambda^{\prime}_{i}=\sigma^{2},\forall i>J$. Furthermore it is shown that [49, Corollary 4.3]
$k_{i}^{2}=\lambda^{\prime}_{i}\left(\lambda^{\prime}_{i}+\mathrm{tr}(C_{{\bm{y}}\mid{\bm{H}}})\right).$ (112)
Hence, we can conclude
$P\left(\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]\leq\varepsilon\right)$
$\geq 1-\frac{4k^{2}(M-J)\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})}{J^{2}N\varepsilon}\sum_{j\leq J}\frac{|s_{\max}|^{2}+2\sigma^{2}}{\lambda_{j}^{2}},$ (113)
with $\lambda_{j}=(\lambda^{\prime}_{j}-\sigma^{2})/J$ for all $j\in\{1,\dots,J\}$ being the eigenvalues of ${\bm{H}}{\bm{H}}^{\mathrm{H}}$ and
$k^{2}=\sigma^{2}\left(\sigma^{2}+\mathrm{tr}(C_{{\bm{y}}\mid{\bm{H}}})\right).$ (114)